OpenStudy (amonoconnor):

Is this the correct sigma notation of the following sequence? -> "(5/2) + (5/4) + (5/6)" In sigma notation: n=3, k=1, the function = 5/(2k) Can anyone verify this analysis? Any and all help is greatly appreciated! :)

jimthompson5910 (jim_thompson5910):

If you mean to write \[\Large \left(\frac{5}{2}\right)+\left(\frac{5}{4}\right)+\left(\frac{5}{6}\right) = \sum_{k=1}^{3} \frac{5}{2k}\] then you are correct.

OpenStudy (amonoconnor):

Yes I did! Thank you very much! :) the Calc feed is dead tonight, and I believe there are a couple more unanswered questions I know my fellow OS students word surely appreciate having answers to. If you feel so inclined. I just know it sucks to sit for hours on end, waiting for clarification, and you were most helpful to me, just now. :)

jimthompson5910 (jim_thompson5910):

I'm glad to be of help, what else do you need checked?

OpenStudy (amonoconnor):

Here's one: http://openstudy.com/study#/updates/533cc4a6e4b0fb082859416d :)

jimthompson5910 (jim_thompson5910):

I'm not sure what the question is asking exactly